sir if, in the electrolysis of aqueous solution of Cucl2, copper is deposited on the negative electrode and clorine gas liberated on the positive electrode at 800K temperature and 1 bar pressure. how much copper and Cl gas will be liberated if 2.0 ampere current is passed for 1 hour????
 

Grade:12

1 Answers

Sunil Kumar FP
9 years ago
CuCl2 -----Cu+Cl2
Cu2+ +2e- ------Cu
2Cl- -------Cl2 +2e-
Thus 2 mole of electron or 2F coloumb of charge is required.
q=it
=2*3600 C
mole of electron=.075 mole
2 mole will deposit 1 mole or 63.5 gm of Cu
Therefore .075 will deposit=.5*.075*63.5 gm =2.38 gm
mole of chlorine=nRT/P =.0375*.0821*800/1=2.4 gm